solve for the ordered pair
2x-y= 5
y= 3x-4
I got x= 9/5 y=-7/5

Answer by checkley71(8403)

2X-Y=5 OR -Y=5-2X OR Y=2X-5 SUBSTITUTING (2X-5) FOR Y IN EQUATION #2 WE GWT
2X-5=3X-4 OR -X=1 OR X=-1 SUBSTITUTING -1 FOR X IN EQUATION 1 WE GET
Y=(3*-1)-4 OR Y=-3-4 OR Y=-7.
DON'T KNOW WHERE YOU WENT WRONG BUT YOU DID!!!!